# ZeroFlow: Overcoming Catastrophic Forgetting is Easier than You Think

### v3

@@ Abstract
Backpropagation provides a generalized configuration for overcoming catastrophic forgetting. Like, SGD and Adam are commonly used for weight updates in continual learning and continual pre-training. In practice, permission to access gradient information is not always granted (the gradient ban), such as black-box APIs, hardware limitations, and non-differentiable systems. To bridge this gap, we introduce the first benchmark ZeroFlow to evaluate gradient-free optimization algorithms for overcoming forgetting. This benchmark examines a suite of forward pass methods across multiple methods, forgetting scenarios, and datasets. We find that forward passes alone are enough to overcome forgetting. Our findings reveal new optimization principles that highlight the potential of forward-pass in mitigating forgetting, managing task conflicts, and reducing memory demands, alongside novel enhancements that further mitigate forgetting with just one forward pass. This work provides essential insights and tools for advancing forward pass methods to overcome forgetting. Code will be available upon publication.
@@ 1 Introduction
Catastrophic forgetting remains one of the major challenges on the path to artificial general intelligence (AGI) (Hadsell et al., 2020; Zhou et al., 2023b), i.e., models tend to forget previously learned tasks when trained on new ones on time-evolving data flow (Feng et al., 2022b). This phenomenon is commonly seen across various tasks, including continual learning (CL) (Wang et al., 2023), fine-tuning of foundation models (FMs) (Mangrulkar et al., 2022; Yuan et al., 2024), and continual pre-training (CPT) (Shi et al., 2024; Zhu et al., 2024b), etc. Among them, optimization algorithms play a crucial role, e.g., SGD has become the default choice during CL (van de Ven et al., 2022), while Adam is frequently seen in fine-tuning FMs (Luo et al., 2023; Zhu et al., 2024a). These optimization algorithms in tandem with various methods (ranging from regularization and rehearsal strategies to architectural changes) rely on gradient information to avoid forgetting (Zhou et al., 2023c; Bian et al., 2024). Nonetheless, in real-world scenarios, gradient information is not always available or computable (i.e., the gradient ban), like, Scenario i: large language models as a service (LLMaaS) and black-box APIs. Scenario ii: hardware systems that do not support principled backpropagation. Scenario iii: AI for science with non-differentiable underlying systems.
@@
To tackle (Q), a natural idea is to use the forward pass method (Hinton, 2022; Baydin et al., 2022; Ren et al., 2022) instead of backpropagation to overcome forgetting. The zeroth-order (ZO) optimization methods (Flaxman et al., 2004; Nesterov & Spokoiny, 2017; Malladi et al., 2023; Ghadimi & Lan, 2013), as representative methods, are well-suited to this issue due to their relaxed information requirements, as they rely only on function values rather than gradients. Under gradient bans, DECL and DFCL (Yang et al., 2024) first attempt to overcome forgetting from a stream of APIs, but they focus on synthetic data level rather than optimization. Therefore, it remains elusive whether benchmark studies using gradient-free methods can mitigate forgetting.
In this work, we explore several Zeroth-order optimization methods on dynamic data Flow (as shown in Figure 1), examining their performance across various forgetting scenarios, model types, and evaluation metrics. Through a detailed analysis, we reveal the overlooked potential of forward passes and various ZO methods in overcoming catastrophic forgetting. This benchmark study offers an easier way to overcome forgetting and helps reveal the pros and cons of these methods in alleviating forgetting. Extended from the gained insights, we introduce a new improvement known as the periodic gradient technique, which further improves ZO optimization for overcoming catastrophic forgetting. Simply put, we can better and more efficiently mitigate forgetting by using just one forward pass.
@@
(ii) Through this benchmark, we uncover overlooked optimization principles and insights into how forward pass can mitigate forgetting. These include the role of forward pass in managing task conflicts and the trade-offs between forgetting and memory efficiency. We proved that catastrophic forgetting can be overcome in an easier way!
(iii) Apart from a comprehensive evaluation of catastrophic forgetting, we introduce three enhancement techniques, which further improve the performance and efficiency of just one forward pass to overcome forgetting.
@@ 2 Literatures
Catastrophic forgetting. Catastrophic forgetting occurs across various tasks, including CL, fine-tuning of FMs, and CPT (Zhou et al., 2023b; Wang et al., 2023; Zhu et al., 2024a; Luo et al., 2023). To mitigate this issue, various methods have been proposed (Rolnick et al., 2019a; Jeeveswaran et al., 2023; Sun et al., 2023b; Li et al., 2024). In CL, methods range from regularization and rehearsal strategies to architectural changes (Kirkpatrick et al., 2017; Bian et al., 2024; Lu et al., 2024). Lately, pre-trained models (PTM) further advanced these methods due to their strong generalization (Yuan et al., 2022; Feng et al., 2022a), as seen in PTM-based CL (Zhou et al., 2024a). All these methods share a common goal: achieving an optimal balance between learning plasticity and memory stability (Wang et al., 2023). In FMs, catastrophic forgetting often arises from overfitting to small fine-tuning datasets during CPT or fine-tuning (Luo et al., 2023; Zhu et al., 2024a). Common techniques to address this include learning rate adjustment, parameter-efficient fine-tuning, mixed data strategies, and instruction tuning (Luo et al., 2023; Hu et al., 2021). Additionally, as foundational models increasingly gain multimodal capabilities, the complexity of catastrophic forgetting also intensifies (Zhu et al., 2024a).
@@
Optimization for catastrophic forgetting. Two broad categories of optimization methods exist for overcoming forgetting, (i) Standard Optimization. SGD and the Adam family are frequently employed to investigate catastrophic forgetting (Cha et al., 2021; Hadsell et al., 2020; Masana et al., 2022). For instance, in CL, various CL methods predominantly utilize the SGD optimizer for standard evaluations (van de Ven et al., 2022; Sun et al., 2023a; Zhou et al., 2024c). In fine-tuning the LLM, the Adam series is commonly used to observe forgetting phenomena (Luo et al., 2023; Zhu et al., 2024a). Some works explored orthogonal spaces with these standard optimizers to alleviate forgetting (Lopez-Paz & Ranzato, 2017; Feng et al., 2022c; Saha et al., 2020), such as OGD (Farajtabar et al., 2020), and GPM (Saha et al., 2020). Moreover, other works (Farajtabar et al., 2020; Chaudhry et al., 2018; Lopez-Paz & Ranzato, 2017) modified the gradients in the standard optimization process to align the learning spaces of new and old tasks, such as Uni-Grad (Li et al., 2024). The core of these efforts (Deng et al., 2021; Shi et al., 2021) is to find an equilibrium between learning and forgetting in optimization. (ii) Sharpness-aware Optimization. This series of methods (He et al., 2019; Foret et al., 2020; Zhong et al., 2022; Zhuang et al., 2022) has gained attention due to the effectiveness of the flat minimum in mitigating forgetting (Li et al., 2024; Kong et al., 2023; Cha et al., 2021; Mehta et al., 2023). Methods such as FS-DPGM (Deng et al., 2021), F2M (Shi et al., 2021), DFGP (Yang et al., 2023), SAM-CL (Tung et al., 2023) overcome forgetting in the flatness areas of different configurations. C-Flat (Bian et al., 2024) proposed a CL-friendly general optimization framework, that holds promise as a baseline optimizer for overcoming forgetting.
@@ 3 Exploring Zeroth-Order Optimization to Overcome Forgetting
3.1 Zeroth-Order Optimization
@@
1) ZO gradient estimation. Randomized Gradient Estimation (RGE (Nesterov & Spokoiny, 2017)) and Coordinate-wise Gradient Estimation (CGE (Berahas et al., 2022)) perturb the model using $\mathcal{\xi}$, which is generated either from a random unknown distribution (in RGE) or by modifying individual coordinates (in CGE), and then observe the changes in the loss function $\mathcal{L}$ after each perturbation, step by step, to provide a reliable gradient estimate. However, due to their reliance on slow single-direction perturbation, these methods are not well-suited for deep learning tasks, as performing a full perturbation in high-dimensional parameter spaces is time-consuming. For instance, the ImageNet have input dimensions of 224×224×3, totaling approximately 0.13 million parameters, and rendering over each dimension iteratively leads to query inefficiency. Standard Simultaneous Perturbation Stochastic Approximation (SPSA) improves efficiency by generating pairs of symmetric vectors and perturbing in multiple directions simultaneously, as follows,
^$\hat{\nabla}L(\theta,\xi;\mathcal{B})=\frac{f(\theta+\epsilon\xi;\mathcal{B})-f(\theta-\epsilon\xi;\mathcal{B})}{2\epsilon}\xi^{-1}.$ (1)
Where $\epsilon$ is a positive scaler and $\xi$ is recommended to follow a symmetric distribution with finite inverse moments (e.g., the Rademacher distribution). The symmetric distribution ensures unbiased exploration of perturbations in both positive and negative directions of parameters at each step. And the finite inverse moments property guarantees that the steps are well-controlled, avoiding excessively large steps due to $\xi^{-1}$ drawn from the distribution (e.g., $\mathbb{E}[1/|\xi|^{p}]$ for some large p), which would otherwise lead to an unstable optimization process. In practical implementations for models with a large number of parameters (e.g., MeZO (Malladi et al., 2023) in large language models (LLMs)), Gaussian noise with zero mean induces substantial perturbations, thereby enhancing exploration across the parameter space and facilitating the escape from local minima. This methodology achieves gradient estimation with only two objective function evaluations, rendering its computational cost independent of input dimensionality. Such computational efficiency has established SPSA as a preferred method for addressing the complexities of high-dimensional deep learning tasks. While increasing $q$ in $q$-SPSA can improve stability in the update direction, setting $q=1$ is sufficient for pretrained LLMs  (Malladi et al., 2023).
@@
3.2 Zeroth-Order Optimization for Catastrophic Forgetting
(a) FO-SGD
(b) ZO-SGD
Figure 3: Trajectory of FO and ZO Optimization during Overcoming Forgetting. The trajectory is taken when using the total loss from both tasks (cyan) and the gradients from each individual task at fixed points during optimization (red and blue). The trends of ZO optimization hold the potential to manage forgetting and learning.
@@ 3.2 Zeroth-Order Optimization for Catastrophic Forgetting
where $N_{context}$ represents the number of contexts encountered so far. In Experience Replay (Rolnick et al., 2019b), both components use classification loss based on their respective data distributions, so ZO gradients can be expressed as ^$\hat{\nabla}\mathcal{L}_{cur}$ and ^$\hat{\nabla}\mathcal{L}_{replay}$ respectively. However, in the emerging generative replay workflows (Shin et al., 2017), Equation 2 may introduce additional loss for the training of generators. In this case, the generator can be trained using standard backpropagation or in conjunction with ZO training without FO gradients.
@@
Table 1: ZeroFlow Evaluation on CIFAR-100, ImageNet-A, CUB and OmniBenchmark. This table compares average accuracy, final accuracy, and forgetting measures of 2 models, and 4 forgetting scenarios. More intuitive trend please see Figure 2. Note that all ZO optimization w/ query budget $q=1$, which means that all evaluations use just one forward pass.
@@
As Figure 3, we visualize and compare the optimization trajectories of ZO and FO methods in the learning-memory trade-off dynamics of continual learning. The remarkable similarity between these trajectories motivates our investigation, suggesting ZO optimization’s potential in effectively managing the balance between learning and forgetting.
Potential. The intrinsic optimization mechanism of ZO exhibits particular promise in continual learning scenarios. Intuitively, ZO perturbs parameters using random or coordinate-wise directional vectors and observes changes in the evaluation function, effectively optimizing within a noisy environment. This approach enables small parameter modifications to yield significant impacts on target objectives, resulting in distinctive gradient estimations compared to FO optimization. Notably, while ZO methods do not explicitly incorporate sharpness regularization terms (Foret et al., 2021), they naturally facilitate the exploration of flat regions in parameter space. The influence of optimizing flat regions with ZO approaches in continual learning can be summarized in two main manifolds: (i) For previous tasks, the noise-induced parameter robustness enhances resilience against perturbations from new task adaptation; (ii) For new tasks, empirical evidence suggests that convergence to flat minima generally leads to lower generalization error.
Risk. Although ZO demonstrates superior generalization abilities, its practical performance is limited by optimization strategies and the complexity of the optimization setting. Despite significant efforts to reduce convergence error, optimizing models from scratch in high-dimensional space remains challenging due to slow convergence speed (proportional to the parameter dimension $d$). For instance, origin CGE-based ZO training for a model with 12k parameters takes 70.32 hours in DeepZero (Chen et al., 2024). Such computational demands render from scratch training impractical for high-dimensional CL models, particularly those employing expansion-based architectures. Consequently, we focus our discussion on leveraging ZO optimization to overcome forgetting within a pre-training context.

### v4

@@ Abstract
Backpropagation provides a generalized configuration for overcoming catastrophic forgetting. Optimizers such as SGD and Adam are commonly used for weight updates in continual learning and continual pre-training. However, access to gradient information is not always feasible in practice due to black-box APIs, hardware constraints, or non-differentiable systems, a challenge we refer to as the gradient bans. To bridge this gap, we introduce ZeroFlow, the first benchmark designed to evaluate gradient-free optimization algorithms for overcoming forgetting. ZeroFlow examines a suite of forward pass-based methods across various algorithms, forgetting scenarios, and datasets. Our results show that forward passes alone can be sufficient to mitigate forgetting. We uncover novel optimization principles that highlight the potential of forward pass-based methods in mitigating forgetting, managing task conflicts, and reducing memory demands. Additionally, we propose new enhancements that further improve forgetting resistance using only forward passes. This work provides essential tools and insights to advance the development of forward-pass-based methods for continual learning.
@@ 1 Introduction
Catastrophic forgetting remains one of the major challenges on the path to artificial general intelligence (AGI) (Hadsell et al., 2020; Zhou et al., 2023b), i.e., models tend to forget previously learned tasks when trained on new ones on time-evolving data flow (Feng et al., 2022b). This phenomenon is commonly seen across various tasks, including continual learning (CL) (Wang et al., 2023), fine-tuning of foundation models (FMs) (Sun et al., 2025; Yuan et al., 2024), and continual pre-training (CPT) (Shi et al., 2024; Zhu et al., 2024b), etc. Among them, optimization algorithms play a crucial role, e.g., SGD has become the default choice during CL (van de Ven et al., 2022), while Adam is frequently seen in fine-tuning FMs (Luo et al., 2023; Zhu et al., 2024a). These optimization algorithms in tandem with various methods (ranging from regularization and rehearsal strategies to architectural changes) rely on gradient information to avoid forgetting (Zhou et al., 2023c; Bian et al., 2024). Nonetheless, in real-world scenarios, gradient information is not always available or computable (i.e., the gradient bans), like, Scenario i: large language models as a service (LLMaaS) and black-box APIs. Scenario ii: hardware systems that do not support principled backpropagation. Scenario iii: AI for science with non-differentiable underlying systems.
@@
To tackle (Q), a natural idea is to use the forward pass-based method (Hinton, 2022; Baydin et al., 2022; Ren et al., 2022) instead of backpropagation to overcome forgetting. The zeroth-order (ZO) optimization methods (Flaxman et al., 2004; Nesterov & Spokoiny, 2017; Malladi et al., 2023; Ghadimi & Lan, 2013), as representative methods, are well-suited to this issue due to their relaxed information requirements, as they rely only on function values rather than gradients. Under gradient bans, DECL and DFCL (Yang et al., 2024) first attempt to overcome forgetting from a stream of APIs, but they focus on synthetic data level rather than optimization. Therefore, it remains elusive whether benchmark studies using gradient-free methods can mitigate forgetting.
In this work, we explore several Zeroth-order optimization methods on dynamic data Flow (as shown in Figure 1), examining their performance across various forgetting scenarios, model types, and evaluation metrics. Through a detailed analysis, we reveal the overlooked potential of forward passes and various ZO methods in overcoming catastrophic forgetting. This benchmark study offers an easier way to overcome forgetting and helps reveal the pros and cons of these methods in alleviating forgetting. Extended from the gained insights, we introduce three new enhancement variants that further improve ZO optimization to overcome catastrophic forgetting. Simply put, we can mitigate forgetting more effectively and efficiently using only forward passes.
@@
(ii) Through this benchmark, we uncover overlooked optimization principles and insights into how forward passes can mitigate forgetting. These include the role of forward passes in managing task conflicts and the trade-offs between forgetting and memory efficiency. We proved that catastrophic forgetting can be overcome in an easier way!
(iii) Apart from a comprehensive evaluation of catastrophic forgetting, we introduce three enhancement techniques, which further improve the performance and efficiency of just forward passes to overcome forgetting.
@@ 2 Literatures
Catastrophic forgetting. Catastrophic forgetting occurs across various tasks, including CL, fine-tuning of FMs, and CPT (Zhou et al., 2023b; Wang et al., 2023; Zhuang et al., 2022a; Luo et al., 2023). To mitigate this issue, various methods have been proposed (Aojun et al., 2025; Jeeveswaran et al., 2023; Sun et al., 2023b; Li et al., 2024). In CL, methods range from regularization and rehearsal strategies to architectural changes (Zhuang et al., 2023; Bian et al., 2024; Lu et al., 2024). Lately, pre-trained models (PTM) further advanced these methods due to their strong generalization (Yuan et al., 2022; Feng et al., 2022a), as seen in PTM-based CL (Zhou et al., 2024a). All these methods share a common goal: achieving an optimal balance between learning plasticity and memory stability (Wang et al., 2023). In FMs, catastrophic forgetting often arises from overfitting to small fine-tuning datasets during CPT or fine-tuning (Luo et al., 2023; Zhu et al., 2024a). Common techniques to address this include learning rate adjustment, parameter-efficient fine-tuning, mixed data strategies, and instruction tuning (Luo et al., 2023; Zhang et al., 2025). Additionally, as foundational models increasingly gain multimodal capabilities, the complexity of catastrophic forgetting also intensifies (Zhao et al., 2024a; Zhu et al., 2024a).
@@
Optimization for catastrophic forgetting. Two broad categories of optimization methods exist for overcoming forgetting, (i) Standard Optimization. SGD and the Adam family are frequently employed to investigate catastrophic forgetting (Hadsell et al., 2020; Masana et al., 2022). For instance, in CL, various CL methods predominantly utilize the SGD optimizer for standard evaluations (van de Ven et al., 2022; Sun et al., 2023a; Zhou et al., 2024c). In fine-tuning the LLM, the Adam series is commonly used to observe forgetting phenomena (Luo et al., 2023; Zhu et al., 2024a). Some works explored orthogonal spaces with these standard optimizers to alleviate forgetting (Lopez-Paz & Ranzato, 2017; Feng et al., 2022c; Saha et al., 2020), such as OGD (Farajtabar et al., 2020), and GPM (Saha et al., 2020). Moreover, other works (Farajtabar et al., 2020; Chaudhry et al., 2018; Lopez-Paz & Ranzato, 2017) modified the gradients in the standard optimization process to align the learning spaces of new and old tasks, such as Uni-Grad (Li et al., 2024). The core of these efforts (Deng et al., 2021; Shi et al., 2021) is to find an equilibrium between learning and forgetting in optimization. (ii) Sharpness-aware Optimization. This series of methods (He et al., 2019; Foret et al., 2020; Zhong et al., 2022; Zhuang et al., 2022b) has gained attention due to the effectiveness of the flat minimum in mitigating forgetting (Li et al., 2024; Kong et al., 2023; Cha et al., 2021; Mehta et al., 2023). Methods such as FS-DPGM (Deng et al., 2021), F2M (Shi et al., 2021), DFGP (Yang et al., 2023), SAM-CL (Tung et al., 2023) overcome forgetting in the flatness areas of different configurations. C-Flat (Bian et al., 2024) proposed a CL-friendly general optimization framework, that holds promise as a baseline optimizer for overcoming forgetting.
@@ 3 Exploring Zeroth-Order Optimization to Overcome Forgetting
C.1 Zeroth-Order Optimization
@@
1) ZO gradient estimation. Randomized Gradient Estimation (RGE (Nesterov & Spokoiny, 2017)) and Coordinate-wise Gradient Estimation (CGE (Berahas et al., 2022)) perturb the model using $\mathcal{\xi}$, which is generated either from a random unknown distribution (in RGE) or by modifying individual coordinates (in CGE), and then observe the changes in the loss function $\mathcal{L}$ after each perturbation, step by step, to provide a reliable gradient estimate. However, due to their reliance on slow single-direction perturbation, these methods are not well-suited for deep learning tasks, as performing a full perturbation in high-dimensional parameter spaces is time-consuming. For instance, typical vision models like ResNet trained on ImageNet have over 25 million parameters. Performing per-dimension perturbations over such a large parameter space renders ZO-based querying highly inefficient. Standard Simultaneous Perturbation Stochastic Approximation (SPSA(Spall, 1992)) improves efficiency by generating pairs of symmetric vectors and perturbing in multiple directions simultaneously, as follows,
^$\hat{\nabla}L(\theta,\xi;\mathcal{B})=\frac{L(\theta+\epsilon\xi;\mathcal{B})-L(\theta-\epsilon\xi;\mathcal{B})}{2\epsilon}\xi^{-1}.$ (1)
Where $\epsilon$ is a positive scaler and $\xi$ is recommended to follow a symmetric distribution with finite inverse moments (e.g., the Rademacher distribution). The symmetric distribution ensures unbiased exploration of perturbations in both positive and negative directions of parameters at each step. And the finite inverse moments property guarantees that the steps are well-controlled, avoiding excessively large steps due to $\xi^{-1}$ drawn from the distribution (e.g., $\mathbb{E}[1/|\xi|^{p}]$ for some large p), which would otherwise lead to an unstable optimization process. In practical implementations for models with a large number of parameters (e.g., MeZO (Malladi et al., 2023) in LLMs (Zhao et al., 2024b)), Gaussian noise with zero mean induces substantial perturbations, thereby enhancing exploration across the parameter space and facilitating the escape from local minima. This methodology achieves gradient estimation with only two objective function evaluations, rendering its computational cost independent of input dimensionality. Such computational efficiency has established SPSA as a preferred method for addressing the complexities of high-dimensional deep learning tasks. While increasing $q$ in $q$-SPSA can improve stability in the update direction, setting $q=1$ is sufficient for pretrained LLMs  (Malladi et al., 2023).
@@
C.2 Zeroth-Order Optimization for Catastrophic Forgetting
(a) FO-Adam
(b) ZO-Adam
Figure 3: Trajectory of FO and ZO Optimization during Overcoming Forgetting. The trajectory is taken when using the total loss from both tasks (cyan) and the gradients from each individual task at fixed points during optimization (red and orange). The trends of ZO optimization hold the potential to manage forgetting and learning.
@@ C.2 Zeroth-Order Optimization for Catastrophic Forgetting
where $N_{context}$ represents the number of contexts encountered so far. In Experience Replay (Rolnick et al., 2019), both components use classification loss based on their respective data distributions, so ZO gradients can be expressed as ^$\hat{\nabla}\mathcal{L}_{cur}$ and ^$\hat{\nabla}\mathcal{L}_{replay}$ respectively. However, in the emerging generative replay workflows (Shin et al., 2017), Equation 2 may introduce additional loss for the training of generators. In this case, the generator can be trained using standard backpropagation or in conjunction with ZO training without FO gradients.
@@
Table 1: ZeroFlow Evaluation on CIFAR-100, ImageNet-A, CUB and OmniBenchmark. This table compares average accuracy, final accuracy, and forgetting measures of 2 models, and 4 forgetting scenarios. More intuitive trend please see Figure 2. All ZO optimizations use a query budget of $q=1$. Bold indicates the best accuracy achieved among ZeroFlow.
@@
As shown in Figure 3, we visualize and compare the optimization trajectories of ZO and FO methods under the learning–memory trade-off dynamics in continual learning. The objective is defined over two-dimensional parameters, with axes specified in Appendix A.2. The striking similarity between the two trajectories highlights the potential of ZO optimization in effectively balancing learning and forgetting, thereby motivating our further investigation.
Potential. The intrinsic optimization mechanism of ZO exhibits particular promise in continual learning scenarios. Intuitively, ZO perturbs parameters using random or coordinate-wise directional vectors and observes changes in the evaluation function, effectively optimizing within a noisy environment. This approach enables small parameter modifications to yield significant impacts on target objectives, resulting in distinctive gradient estimations compared to FO optimization. Notably, while ZO methods do not explicitly incorporate sharpness regularization terms, they naturally facilitate the exploration of flat regions in parameter space. The influence of optimizing flat regions with ZO approaches in continual learning can be summarized in two main manifolds: (i) For previous tasks, the noise-induced parameter robustness enhances resilience against perturbations from new task adaptation; (ii) For new tasks, empirical evidence suggests that convergence to flat minima generally leads to lower generalization error.
Risk. Although ZO demonstrates superior generalization abilities, its practical performance is limited by optimization strategies and the complexity of the optimization setting. Despite significant efforts to reduce convergence error, optimizing models from scratch in high-dimensional space remains challenging due to slow convergence speed (proportional to the parameter dimension $d$). For instance, origin CGE-based ZO training for a model with 12k parameters takes 70.32 hours in DeepZero (Chen et al., 2023). Such computational demands render from scratch training impractical for high-dimensional CL models, particularly those employing expansion-based architectures. Consequently, we focus our discussion on leveraging ZO optimization to overcome forgetting within a pre-training context.